A pair of dice his tossed twice. Find the probability that both roles give a sum of five

I got 1/9 and it was wrong
Answer #1

here 1/9 is probabilty of a sum of 5 on single roll of pair of dice

while for 2 rolls ;; probability that sum is 5 on both rolls of pair of dice=(1/9)*(1/9)=1/81

( please revert)